shows a piece of block is 15 cm long, 7 cm wide and 3 cm thick is floating with one-third of its volume above the water level. Calculate the buoyant force exerted on the block.


1
Expert's answer
2022-01-24T12:02:24-0500

Given:

"V=\rm (15\:cm)\times(7\:cm)\times (3\:cm)=315\; cm^3"

"V_s=2/3V=\rm 210\; cm^3=2.1*10^{-4}\:m^3"

"\rho=\rm 1000\:kg/m^3"


The buoyant force

"F_b=g\rho V_s=9.8*1000*2.1*10^{-4}=2.06\: \rm N"